Solve the inequality, state the solution set using interval notation
1/2x^2 is greater than/or equal to 4-x

Answer by smartdude17(30)

I don't know what you mean by interval notation, but this is what I got:
1/2x^2 >/= 4-x G
1/2x^2 >/= 4-x]2 I(CD)
x^2 >/= 8-2x S
x^2+2x-8= 0 I,change to =
(x+4)(x-2)= 0 FoiL (you can also do the QF)
x+4=0, x-2=0 Set all terms to zero
x=-4, x=2 Isolate,S
And here are your answers. If you have to graph it, use -b/2a.